Please clarify me with the difference of Line,single,double.triple,surface integrals.


Please respond me soon


Also tell me how good is the college IEM Salt Lake Kolkata and how the future is in this college and abouts its placement.

Grade:12

1 Answers

AVISHKAR JADON
13 years ago

HI..   ACTUALLY LINE INTEGRAL IS USED TO INTEGRATE SOMETHING WHICH IS LINEAR OR WHICH EXISTS IN 1D FOR EX.   ∫B.dl =Μ0i   .. SIMILARLY FOR INTEGRATING SOMETHING WHICH IS 2D LIKE TO FIND AREA WE USE SURFACE INTEGRALS AND FOR VOLUME OR 3D WE   USE  CUBIC INTEGRALS OR CLOSED INTEGRALS .  DOUBLE OR TRIPLE INTEGRAL IS NOTHING BUT ONLY  USED TO DO INTEGRATION MORE THAN ONE TIME .

FOR EX..∫∫∫x+1 =∫∫(x2/2+x)   =∫x3/6+x2/2 =x4/24+x3/6..

by.... waiting for more......